# The computational power of discrete chemical reaction networks with bounded executions

### v3

@@ Abstract
However, if no initial leader is allowed, all species vote, and the CRN is “non-collapsing” (does not shrink from initially large to final $O(1)$ size configurations), then execution bounded CRNs are severely limited, able to compute only eventually constant predicates. A key tool is a characterization of execution bounded CRNs as precisely those with a nonnegative linear potential function that is strictly decreased by every reaction [6].
@@ 1 Introduction
Chemical reaction networks (CRNs) are a fundamental tool for understanding and designing molecular systems. By abstracting chemical reactions into a set of finite, rule-based transformations, CRNs allow us to model the behavior of complex chemical systems. For instance, the CRN with a single reaction $2X\rightarrow Y$, produces one $Y$ every time two $X$ molecules randomly react together, effectively calculating the function $f(x)=\lfloor x/2\rfloor$ if the initial count of $X$ is interpreted as the input and the eventual count of $Y$ as the output. A commonly studied special case of CRNs is the population protocol model of distributed computing [3], in which each reaction has exactly two reactants and two products, e.g., $A+B\to C+D$. This model assumes idealized conditions where reactions can proceed indefinitely, constrained only by the availability of reactants in the well-mixed solution.
Precisely the semilinear predicates $\phi:\mathbb{N}^{d}\to\{0,1\}$ [1] and functions $f:\mathbb{N}^{d}\to\mathbb{N}$ [5] can be computed stably, roughly meaning that the output is correct no matter the order in which reactions happen. In population protocols or other CRNs with a finite reachable configuration space, this means that the output is correct with probability 1 under a stochastic scheduler that picks the next molecules to react at random. However, existing constructions to compute semilinear predicates and functions use CRNs with unbounded executions, meaning that it is possible to execute infinitely many reactions from the initial configuration. CRNs with bounded executions have several advantages. With an absolute guarantee on how many reactions will happen before the CRN terminates, wet-lab implementations need only supply a bounded amount of fuel to power the reactions. Such CRNs are simpler to reason about: each reaction brings it “closer” to the answer. They also lead to a simpler definition of stable computation than is typically employed: an execution bounded CRN stably computes a predicate/function if it gets the correct answer after sufficiently many reactions.
To study this topic, we study networks that must eventually reach a configuration where no further reactions can occur, regardless of the sequence of reactions executed. This restriction is nontrivial because the techniques of [5, 7] rely on reversible reactions (leading to unbounded executions) catalyzed by species we expect to be depleted once a computational step has terminated. This trick seems to add computational power to our system by undoing certain reactions as long as a specific species is present. Consider the following CRN computing $f(x_{1},x_{2},x_{3})=\min\left(x_{1}-x_{2},x_{3}\right)$. The input values $x_{i}$ are given by the counts of $X_{i}$, and the output by the count of $Z$ molecules in the stable state:
@@
Reactions (1) and (2) compute $x_{1}-x_{2}$, storing the result in the count of $Y$. Next, reaction (3) can be applied exactly $\min(y,x_{3})$ times. But since the order of reactions is a stochastic process, we might consume copies of $Y$ in (3), before all of $x_{2}$ is subtracted from it. Therefore, we add reaction (4), using $X_{2}$ as a catalyst to undo reaction (3) as long as copies of $X_{2}$ are present, indicating that the first step of computation has not terminated. However, this means the above CRN does not have bounded exectutions, since reactions (3) and (4) can be alternated in an infinite execution. A similar technique is used in [5], where semilinear sets are understood as a finite union of linear sets, shown to be computable in parallel by CRNs. A reversible, catalyzed reaction finally converts the output of one of the CRNs to the global output. Among other questions, we explore how the constructions of [5] and [7] can be modified to provide equal computational power while guaranteeing bounded execution.
The paper is organized as follows. Section 3 defines execution boundedness (Definition 3.1). We introduce alternative characterizations of the class for use in later proofs, such as the lack of self-covering execution paths. Section 4 and 5 contain the main positive results of the paper and provide the concrete constructions used to decide semilinear sets and functions using execution bounded CRNs whose initial configurations contain a single leader. Section 6 discusses the limitations of execution bounded CRNs, introducing the concept of a “linear potential function” as a core characterization of these systems. We demonstrate that entirely execution bounded CRNs that are leaderless and non-collapsing (such as all population protocols), can only stably decide trivial semilinear predicates: the eventually constant predicates (Definition 6.11).
@@ 2 Preliminaries
We use established notation from [5, 7] and stable computation definitions from [3] for (discrete) chemical reaction networks.
@@ 2.2 Chemical Reaction Networks
A chemical reaction network (CRN) is a pair $\mathcal{C}=(\Lambda,R)$, where $\Lambda$ is a finite set of chemical species, and $R$ is a finite set of reactions over $\Lambda$, where each reaction is a pair $(\mathbf{r},\mathbf{p})\in\mathbb{N}^{\Lambda}\times\mathbb{N}^{\Lambda}$ indicating the reactants $\mathbf{r}$ and products $\mathbf{p}$. A population protocol [1] is a CRN in which all reactions $(\mathbf{r},\mathbf{p})$ obey $\|\mathbf{r}\|=\|\mathbf{p}\|=2$. (Note that CRNs, including population protocols, do not assume any underlying “communication graph” and model a well-mixed system in which each equal-sized of molecules is as likely to collide and react as any other.) We write reactions such as $A+2B\mathop{\rightarrow}\limits A+3C$ to represent the reaction ($(\{A,2B\},\{A,3C\})$. A configuration $\mathbf{c}\in\mathbb{N}^{\Lambda}$ of a CRN assigns integer counts to every species $S\in\Lambda$. When convenient, we use the notation $\left\{n_{1}S_{1},n_{2}S_{2},\ldots,n_{k}S_{k}\right\}$ to describe a configuration $\mathbf{c}$ with $n_{i}\in\mathbb{N}$ copies of species $S_{i}$, i.e., $\mathbf{c}(S_{i})=n_{i}$, and any species that is not listed is assumed to have a zero count. If some configuration $\mathbf{c}$ is understood from context, for a species $S$, we write #$\#S$ to denote $\mathbf{c}(S).$ A reaction $(\mathbf{r},\mathbf{p})$ is said to be applicable in configuration $\mathbf{c}$ if $\mathbf{r}\leqq\mathbf{c}$. If the reaction $(\mathbf{r},\mathbf{p})$ is applicable, applying it results in configuration $\mathbf{c}^{\prime}=\mathbf{c}-\mathbf{r}+\mathbf{p}$, and we write $\mathbf{c}\rightarrow\mathbf{c}^{\prime}$.
@@
[…]
@@
[…]
submission,final $\mathbf{M}=\begin{pmatrix}-1&1\\ 1&-2\\ 2&0\end{pmatrix}.$
@@ 2.4 Time model
The following model of stochastic chemical kinetics is widely used in quantitative biology and other fields dealing with chemical reactions between species present in small counts [10]. It ascribes probabilities to execution sequences, and also defines the time of reactions, allowing us to study the computational complexity of the CRN computation in Sections 4 and 5. If the volume is defined to be the total number of molecules, then the time model is essentially equivalent to the notion of parallel time studied in population protocols [2]. In this paper, the rate constants of all reactions are $1$, and we define the kinetic model with this assumption. A reaction is unimolecular if it has one reactant and bimolecular if it has two reactants. We use no higher-order reactions in this paper.
@@
The kinetic model is based on the physical assumption of well-mixedness that is valid in a dilute solution. Thus, we assume the finite density constraint, which stipulates that a volume required to execute a CRN must be proportional to the maximum molecular count obtained during execution [17]. In other words, the total concentration (molecular count per volume) is bounded. This realistically constrains the speed of the computation achievable by CRNs.
@@ 2.5 Semilinear sets, predicates, functions
The following is a known characterization of the computational power of CRNs [3, 4].
Theorem 2.3 ([3, 4]).
@@ 3 Execution bounded chemical reaction networks
[…]
@@
This is a distinct concept from the notion of “bounded” CRNs studied by Rackoff [16] (studied under the equivalent formalism of vector addition systems). That paper defines a CRN to be bounded from a configuration $\mathbf{x}$ if $|\mathsf{reach}(\mathbf{x})|$ is finite (and shows that the decision problem of determining whether this is true is $\mathsf{EXPSPACE}$-complete.) We use the term execution bounded to avoid confusion with this concept.
{toappendix}
We first observe that being execution bounded from $\mathbf{x}$ implies a slightly stronger condition: there is a uniform upper bound on the length of all executions from $\mathbf{x}$111 In other words, this rules out the possibility that, although all executions from $\mathbf{x}$ are finite, there are infinitely many of them $\mathcal{E}_{1},\mathcal{E}_{2},\dots$, each finite but longer than the previous.
{observation}
A CRN is execution bounded from $\mathbf{x}$ if and only if there is a constant $N\in\mathbb{N}$ such that all executions from $\mathbf{x}$ have length at most $N$. Equivalently, there are finitely many executions from $\mathbf{x}$, all of which are finite.
Proof 3.2.
We use Kőnig’s lemma to show that in the absence of an infinite path, the number of all possible paths must be finite, which directly implies a global bound on the length of all executions. We represent the set of all executions for $\mathcal{C}$ as a tree where each edge represents a single reaction applied and each node stores the complete execution sequence starting from configuration $\mathbf{x}$. Note that this construction is slightly different from a more straightforward graph with the reachable states as nodes, which would not give us a tree, since the same state can be reached by different executions. Formally, we generate the tree as follows: $T_{\mathcal{C}}(\mathbf{x})=(V,E)$ where $V\triangleq\{\mathcal{E}\in\{\mathbb{N}^{\Lambda}\}^{*}\mid\mathcal{E}\text{ is a valid execution sequence starting from }\mathbf{x}\}$, $E\triangleq\{(\mathcal{E}_{1},\mathcal{E}_{2})\mid\mathcal{E}_{1}\preceq\mathcal{E}_{2}\land|\mathcal{E}_{2}|=|\mathcal{E}_{1}|+1$}. In other words, all the executions from $\mathbf{x}$ of length $d$ are the nodes at depth $d$ of this tree. One can think of the nodes as being labeled by configurations rather than executions (specifically the final configuration of the execution, with the tree rooted at $\mathbf{x}$), but the same configuration can label multiple nodes if it can be reached from $\mathbf{x}$ via different executions. In this case the children of a configuration are those that are reachable from it by applying a single reaction.
This tree is finitely branching, as we can only choose from a finite number of reactions at any node. By definition of execution bounded, there is no execution sequence with an infinite length. Due to the bijection between paths in $T_{\mathcal{C}}(\mathbf{x})$ and executions possible in $\mathcal{C}$, there is no infinite path in the tree. By Kőnig’s Lemma, the tree has a finite number of nodes, guaranteeing a single bound $N$ (the depth of the tree) on the length of every execution.
The next lemma characterizes execution boundedness as equivalent to having a finite reachable state space with no cycles.
@@
A CRN is execution bounded from $\mathbf{x}$ if and only if $\mathsf{reach}(\mathbf{x})=\{\mathbf{y}\mid\mathbf{x}\Rightarrow\mathbf{y}\}$ is finite and, for all $\mathbf{y}\in\mathsf{reach}(\mathbf{x})$, $\mathbf{y}\not\Rightarrow\mathbf{y}$ except by the zero-length execution.
@@
Every configuration reachable from $\mathbf{x}$ is reached through some execution contained in $T_{\mathcal{C}}(\mathbf{x})$ as a node and there exists only a finite number of them (Section 3). Multiple unique executions can produce the same configuration but one execution cannot produce multiple configurations. Thus, there exists a surjection from the nodes of $T_{\mathcal{C}}(\mathbf{x})$ into $\mathsf{reach}(\mathbf{x})$ and $\mathsf{reach}(\mathbf{x})$ must also be finite. For the second part of the condition, we prove its contrapositive and assume there exists $\mathbf{y}\in\mathsf{reach}(\mathbf{x})$, $\mathbf{y}\Rightarrow_{P}\mathbf{y}\land|P|>0$. Let $P=(\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{n})$. It holds that $\mathbf{p}_{1}=\mathbf{p}_{n}$ and $\mathbf{p}_{n-1}\Rightarrow\mathbf{p}_{1}$. We can construct an infinite-length execution $P^{\prime}=(\mathbf{x},\dots,\mathbf{p}_{1},\dots,\mathbf{p}_{n-1},\mathbf{p}_{1},\dots)$, which must also be a valid under the reactions of $\mathcal{C}$, making $\mathcal{C}$ execution unbounded from $\mathbf{x}$.
If $\mathsf{reach}(\mathbf{x})$ is finite and contains no such $\mathbf{y}$, then we can construct a finite, directed, acyclic graph $G_{\mathcal{C}}(\mathbf{x})=(V,E)$ where $V=\mathsf{reach}(\mathbf{x})$, $E=\{(\mathbf{x},\mathbf{y})\mid\mathbf{x}\Rightarrow_{P}\mathbf{y}\land|P|>0\}$. The longest path in the graph has length of at most $|\mathsf{reach}(\mathbf{x})|-1$. A bijection exists between paths in $G_{\mathcal{C}}(\mathbf{x})$ and executions possible in $\mathcal{C}$ starting from $\mathbf{x}$. We set $n=|\mathsf{reach}(\mathbf{x})|$ satisfying that each execution has length of at most $n$, making $\mathcal{C}$ execution bounded.
The following result is used frequently in impossibility proofs for CRNs and population protocols, and it will help us prove another characterization of execution bounded CRNs in Section 3.
Lemma 3.5.
(Dickson’s Lemma) For every infinite sequence of nonnegative integer vectors $\mathbf{x}_{1},\mathbf{x}_{2},\dots\in\mathbb{N}^{k}$, there are $i<j$ such that $\mathbf{x}_{i}\leqq\mathbf{x}_{j}$.
[…]
Definition 3.6.
A execution $\mathcal{E}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots)$ is self-covering if for some $i<j$, $\mathbf{x}_{i}\leqq\mathbf{x}_{j}$. It is strictly self-covering if $\mathbf{x}_{i}\leq\mathbf{x}_{j}$. We also refer to these as (strict) self-covering paths.222 Rackoff [16] uses the term “self-covering” to mean what we call strictly self-covering here, and points out that Karp and Miller [13] showed that $|\mathsf{reach}(\mathbf{x})|$ is infinite if and only if there is a strictly self-covering path from $\mathbf{x}$. The distinction between these concepts is illustrated by the CRN $A\mathop{\rightleftharpoons}\limits B$. From any configuration $\mathbf{x}$, $\mathsf{reach}(\mathbf{x})$ is finite ($|\mathsf{reach}(\mathbf{x})|=\mathbf{x}(A)+\mathbf{x}(B)+1$), and there is no strict self-covering path. However, from (say) $\{A\}$, there is a (nonstrict) self-covering path $\{A\}\Rightarrow\{B\}\Rightarrow\{A\}$, and by repeating, this CRN has an infinite cycling execution within its finite configuration space $\mathsf{reach}(\{A\})=\{\{A\},\{B\}\}$.
{lemmarep}
A CRN is execution bounded from $\mathbf{x}$ if and only if there is no self-covering path from $\mathbf{x}$.
Proof 3.7.

### v4

@@ Abstract
However, if no initial leader is allowed, all species vote, and the CRN is “non-collapsing” (does not shrink from initially large to final $O(1)$ size configurations), then execution bounded CRNs are severely limited, able to compute only eventually constant predicates. A key tool is a characterization of execution bounded CRNs as precisely those with a nonnegative linear potential function that is strictly decreased by every reaction [czerner2024fast].
@@ 1 Introduction
Chemical reaction networks (CRNs) are a fundamental tool for understanding and designing molecular systems. By abstracting chemical reactions into a set of finite, rule-based transformations, CRNs allow us to model the behavior of complex chemical systems. For instance, the CRN with a single reaction $2X\rightarrow Y$, produces one $Y$ every time two $X$ molecules randomly react together, effectively calculating the function $f(x)=\lfloor x/2\rfloor$ if the initial count of $X$ is interpreted as the input and the eventual count of $Y$ as the output. A commonly studied special case of CRNs is the population protocol model of distributed computing [angluin2007computational], in which each reaction has exactly two reactants and two products, e.g., $A+B\to C+D$. This model assumes idealized conditions where reactions can proceed indefinitely, constrained only by the availability of reactants in the well-mixed solution.
Precisely the semilinear predicates $\phi:\mathbb{N}^{d}\to\{0,1\}$ [angluin2004computation] and functions $f:\mathbb{N}^{d}\to\mathbb{N}$ [Chen2012DeterministicFunction] can be computed stably, roughly meaning that the output is correct no matter the order in which reactions happen. In population protocols or other CRNs with a finite reachable configuration space, this means that the output is correct with probability 1 under a stochastic scheduler that picks the next molecules to react at random. However, existing constructions to compute semilinear predicates and functions use CRNs with unbounded executions, meaning that it is possible to execute infinitely many reactions from the initial configuration. CRNs with bounded executions have several advantages. With an absolute guarantee on how many reactions will happen before the CRN terminates, wet-lab implementations need only supply a bounded amount of fuel to power the reactions. Such CRNs are simpler to reason about: each reaction brings it “closer” to the answer. They also lead to a simpler definition of stable computation than is typically employed: an execution bounded CRN stably computes a predicate/function if it gets the correct answer after sufficiently many reactions.
To study this topic, we study networks that must eventually reach a configuration where no further reactions can occur, regardless of the sequence of reactions executed. This restriction is nontrivial because the techniques of [Chen2012DeterministicFunction, Doty2013Leaderless] rely on reversible reactions (leading to unbounded executions) catalyzed by species we expect to be depleted once a computational step has terminated. This trick seems to add computational power to our system by undoing certain reactions as long as a specific species is present. Consider the following CRN computing $f(x_{1},x_{2},x_{3})=\min\left(x_{1}-x_{2},x_{3}\right)$. The input values $x_{i}$ are given by the counts of $X_{i}$, and the output by the count of $Z$ molecules in the stable state:
@@
Reactions (1) and (2) compute $x_{1}-x_{2}$, storing the result in the count of $Y$. Next, reaction (3) can be applied exactly $\min(y,x_{3})$ times. But since the order of reactions is a stochastic process, we might consume copies of $Y$ in (3), before all of $x_{2}$ is subtracted from it. Therefore, we add reaction (4), using $X_{2}$ as a catalyst to undo reaction (3) as long as copies of $X_{2}$ are present, indicating that the first step of computation has not terminated. However, this means the above CRN does not have bounded exectutions, since reactions (3) and (4) can be alternated in an infinite execution. A similar technique is used in [Chen2012DeterministicFunction], where semilinear sets are understood as a finite union of linear sets, shown to be computable in parallel by CRNs. A reversible, catalyzed reaction finally converts the output of one of the CRNs to the global output. Among other questions, we explore how the constructions of [Chen2012DeterministicFunction] and [Doty2013Leaderless] can be modified to provide equal computational power while guaranteeing bounded execution.
The paper is organized as follows. Section˜3 defines execution boundedness (Definition˜3.1). We introduce alternative characterizations of the class for use in later proofs, such as the lack of self-covering execution paths. Section˜4 and 5 contain the main positive results of the paper and provide the concrete constructions used to decide semilinear sets and functions using execution bounded CRNs whose initial configurations contain a single leader. Section˜6 discusses the limitations of execution bounded CRNs, introducing the concept of a “linear potential function” as a core characterization of these systems. We demonstrate that entirely execution bounded CRNs that are leaderless and non-collapsing (such as all population protocols), can only stably decide trivial semilinear predicates: the eventually constant predicates (Definition˜6.9).
@@ 2 Preliminaries
We use established notation from [Chen2012DeterministicFunction, Doty2013Leaderless] and stable computation definitions from [angluin2007computational] for (discrete) chemical reaction networks.
@@ 2.2 Chemical Reaction Networks
A chemical reaction network (CRN) is a pair $\mathcal{C}=(\Lambda,R)$, where $\Lambda$ is a finite set of chemical species, and $R$ is a finite set of reactions over $\Lambda$, where each reaction is a pair $(\mathbf{r},\mathbf{p})\in\mathbb{N}^{\Lambda}\times\mathbb{N}^{\Lambda}$ indicating the reactants $\mathbf{r}$ and products $\mathbf{p}$. A population protocol [angluin2004computation] is a CRN in which all reactions $(\mathbf{r},\mathbf{p})$ obey $\|\mathbf{r}\|=\|\mathbf{p}\|=2$. (Note that CRNs, including population protocols, do not assume any underlying “communication graph” and model a well-mixed system in which each equal-sized of molecules is as likely to collide and react as any other.) We write reactions such as $A+2B\mathop{\rightarrow}\limits A+3C$ to represent the reaction $(\{A,2B\},\{A,3C\})$. A configuration $\mathbf{c}\in\mathbb{N}^{\Lambda}$ of a CRN assigns integer counts to every species $S\in\Lambda$. When convenient, we use the notation $\left\{n_{1}S_{1},n_{2}S_{2},\ldots,n_{k}S_{k}\right\}$ to describe a configuration $\mathbf{c}$ with $n_{i}\in\mathbb{N}$ copies of species $S_{i}$, i.e., $\mathbf{c}(S_{i})=n_{i}$, and any species that is not listed is assumed to have a zero count. If some configuration $\mathbf{c}$ is understood from context, for a species $S$, we write #S\#S to denote $\mathbf{c}(S).$ A reaction $(\mathbf{r},\mathbf{p})$ is said to be applicable in configuration $\mathbf{c}$ if $\mathbf{r}\leqq\mathbf{c}$. If the reaction $(\mathbf{r},\mathbf{p})$ is applicable, applying it results in configuration $\mathbf{c}^{\prime}=\mathbf{c}-\mathbf{r}+\mathbf{p}$, and we write $\mathbf{c}\rightarrow\mathbf{c}^{\prime}$.
@@
For a CRN $\mathcal{C}=(\Lambda,R)$ where $|\Lambda|=n$ and $|R|=m$, define the $n\times m$ stoichiometric matrix $\mathbf{M}$ of $\mathcal{C}$ as follows. The species are ordered $S_{1},\dots,S_{n}$, and the reactions are ordered $(\mathbf{r}_{1},\mathbf{p}_{1}),\dots,(\mathbf{r}_{m},\mathbf{p}_{m})$, and $\mathbf{M}_{ij}=\mathbf{p}_{j}(S_{i})-\mathbf{r}_{j}(S_{i})$. In other words, $\mathbf{M}_{ij}$ is the net amount of $S_{i}$ produced when executing the $j$’th reaction. For instance, if the CRN has two reactions $S_{1}\mathop{\rightarrow}\limits S_{2}+2S_{3}$ and $3S_{2}+S_{3}\mathop{\rightarrow}\limits S_{1}+S_{2}+S_{3}$, then
@@ 2.4 Time model
The following model of stochastic chemical kinetics is widely used in quantitative biology and other fields dealing with chemical reactions between species present in small counts [Gillespie77]. It ascribes probabilities to execution sequences, and also defines the time of reactions, allowing us to study the computational complexity of the CRN computation in Sections˜4 and 5. If the volume is defined to be the total number of molecules, then the time model is essentially equivalent to the notion of parallel time studied in population protocols [AngluinAE2008Fast]. In this paper, the rate constants of all reactions are $1$, and we define the kinetic model with this assumption. A reaction is unimolecular if it has one reactant and bimolecular if it has two reactants. We use no higher-order reactions in this paper.
@@
The kinetic model is based on the physical assumption of well-mixedness that is valid in a dilute solution. Thus, we assume the finite density constraint, which stipulates that a volume required to execute a CRN must be proportional to the maximum molecular count obtained during execution [SolCooWinBru08]. In other words, the total concentration (molecular count per volume) is bounded. This realistically constrains the speed of the computation achievable by CRNs.
@@ 2.5 Semilinear sets, predicates, functions
The following is a known characterization of the computational power of CRNs [angluin2007computational, chen2023rate].
Theorem 2.3 ([angluin2007computational, chen2023rate]).
@@ 3 Execution bounded chemical reaction networks
In this section, we define execution bounded CRNs and state some alternate characterizations of the definition.
@@
This is a distinct concept from the notion of “bounded” CRNs studied by Rackoff [Rackoff1978CoveringBoundedness] (studied under the equivalent formalism of vector addition systems). That paper defines a CRN to be bounded from a configuration $\mathbf{x}$ if $|\mathsf{reach}(\mathbf{x})|$ is finite (and shows that the decision problem of determining whether this is true is $\mathsf{EXPSPACE}$-complete.) We use the term execution bounded to avoid confusion with this concept.
We first observe an equivalent characterization of execution bounded that will be useful in the negative results of Section˜6.
Definition 3.2.
A execution $\mathcal{E}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots)$ is self-covering if for some $i<j$, $\mathbf{x}_{i}\leqq\mathbf{x}_{j}$. It is strictly self-covering if $\mathbf{x}_{i}\leq\mathbf{x}_{j}$. We also refer to these as (strict) self-covering paths. [^1] Rackoff [Rackoff1978CoveringBoundedness] uses the term “self-covering” to mean what we call strictly self-covering here, and points out that Karp and Miller [karp1969parallel] showed that $|\mathsf{reach}(\mathbf{x})|$ is infinite if and only if there is a strictly self-covering path from $\mathbf{x}$. The distinction between these concepts is illustrated by the CRN $A\mathop{\rightleftharpoons}\limits B$. From any configuration $\mathbf{x}$, $\mathsf{reach}(\mathbf{x})$ is finite ($|\mathsf{reach}(\mathbf{x})|=\mathbf{x}(A)+\mathbf{x}(B)+1$), and there is no strict self-covering path. However, from (say) $\{A\}$, there is a (nonstrict) self-covering path $\{A\}\Rightarrow\{B\}\Rightarrow\{A\}$, and by repeating, this CRN has an infinite cycling execution within its finite configuration space $\mathsf{reach}(\{A\})=\{\{A\},\{B\}\}$.
@@
lem:execution-bdd-iff-no-self-covering-path A CRN is execution bounded from $\mathbf{x}$ if and only if there is no self-covering path from $\mathbf{x}$.